Shift report 20010927 00:00-08:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Thu Sep 27 2001 - 08:13:37 EDT

  • Next message: Jens Jørgen Gaardhøje: "Detector workshop RHIC - and BRAHMS"

    ------------------------------------------------------------
    Date: Thu Sep 27 08:13:37 2001, Shift: 00:00-08:00
    Supervisor: J.I.Jordre, with: NN
    Log book#9 pages 54 - 55
    
    Shift Summary:
    ------------------------------------------------------------
    1. Beam from 03:45 to 06:22. One "Cryogenic PLC failure which took
       about 1.5 hours to fix. Now at 7:45 the beam (read: a failed store)
       was dumped for cryo ring entry.
    
    2. Run 4980	Data run with 173.8k trigger 1 events
       Run 4981	TPC pedestal run
       Run 4982	DVM run
       Run 4983	Pedestal run without TPCs
       (FS @ 3 deg., A pol., 1/3 field, MRS @ 90 deg., B pol., I=700A)
    
    3. TOFW Bottom Ch. 25 not good => Dana is attacking the problem. 
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Thu Sep 27 2001 - 08:13:56 EDT